# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ I.4.6 Space-time-optimal resource estimation
Here, we have defined the hybrid (intra- and inter-modular) time required to prepare the subgraph for preparation step $i$ as $t^{\text{prep}}_{i}=8d(N_{i}^{\text{intra-comms}}t+N_{i}^{\text{cross-fridges}}t_{\text{inter}})$. Where $N_{i}^{\text{intra-comms}}=\sum_{j\in L_{i,j}^{\text{prep}}}\lfloor\frac{\mathbb{L}_{i}^{\text{prep}}}{\sum_{k\in L_{i,j}^{\text{prep}}}\lfloor d^{\text{max}}_{i,j,k}/\tilde{n}_{\text{logical}}\rfloor}\rfloor~{}\text{for}~{}n^{\text{modules}}_{\text{per-leg}}>1$ is the count of intra-module operations, while $N_{i}^{\text{cross-fridge}}=\sum_{j\in L_{i}^{\text{prep}}}\lceil\frac{\sum_{k\in L_{i,j}^{\text{prep}}}\lfloor d^{\text{max}}_{i,j,k}/\tilde{n}_{\text{logical}}\rfloor}{n_{\text{inter-pipes}}}\rceil~{}\text{for}~{}n^{\text{modules}}_{\text{per-leg}}>1$ represents the number of cross-fridge operations in the preparation sub-schedule $i$ (only in the vertical direction along a ladder leg). We assume that $t_{\text{inter}}\gg t$ is the characteristic physical timescale for all inter-module communications. Moreover, we define $d^{\text{max}}_{i,j,k}$ as the maximum distance between any pairs of graph nodes in the tuple $k$ of sub-list $j$ of preparation sub-schedule $i$ (see [44] for details), and $n_{\text{inter-pipes}}$ is the number of inter-modular pipes connecting each pair of modules in the macro-architecture.
@@
$t^{\text{distill-delay}}_{\text{i}}=\begin{cases}8tC&\lceil\frac{L^{\text{max-seq}}_{\text{consump},i}-n^{T-\text{per-fridge}}_{i}}{n_{\text{distilleries}}}\rceil,\\ &\text{if }L^{\text{max-seq}}_{\text{consump},i}>n^{T-\text{per-fridge}}_{i}\\ 0,&\text{otherwise}\end{cases}$
@@
Similarly, for step $i$, an idle delay must be added if the preparation for the next subgraph $i+1$ is not yet complete (which runs parallel across the interleaving fridges in the other leg of the ladder). One can establish an upper bound for the individual time related to the intra-modular node consumption of step $i$ as $t^{\text{consump-intra}}_{i}=8td(\lceil\frac{n^{\text{max}-T}_{i}}{n_{\text{distilleries}}}\rceil+L_{\varepsilon}\lceil\frac{n^{\text{max}-Rz}_{i}}{n_{\text{distilleries}}}\rceil)$. Therefore, the preparation delay for step $i\in\{0,\dots,n_{\text{widgets}}-2\}$ becomes
@@
Lastly, we need to account for the total time required on the actual FTQC to stitch the output nodes of the graph at step $i$, located along one leg of the ladder, to the input nodes of the graph at step $i+1$, situated along the other leg of the ladder, for $i\neq n_{\text{widget}}-1$. We assume we can perform the hand-over process on-the-fly by teleporting data-qubit Bell pairs across the modules in a horizontal direction, which involves adding logical $CNOT$ or $CZ$ gates (or, for smaller problems, some equivalent $SWAP$ operations). We must add a cost if FTQC needs to teleport the graph nodes vertically across the same leg (infeasible through $O(1)$ operations). Therefore, an upper bound for this quantity can be expressed as
$\begin{split}t^{\text{total}}_{\text{handover-inter}}=8t_{\text{inter}}d\sum_{i\in\{0,\dots,n_{\text{widget}}-2\}}\lceil\frac{N^{\text{cross-module-IO}}_{i}}{n_{\text{inter-pipes}}}\rceil,\end{split}$ (21)
where we define the total number of cross-fridge communications needed to hand over output nodes of graph $i$ to $i+1$ as $N^{\text{cross-module-IO}}_{i}=\sum_{\tilde{o}\in S^{i}_{\text{output}},\tilde{i}\in S^{i+1}_{\text{input}}}n^{i}_{\tilde{o},\tilde{i}}$; here, $n^{i}_{\tilde{o},\tilde{i}}$ is the number of crossings necessary to hand over node $\tilde{o}$ of graph $i$ to node $\tilde{i}$ of graph $i+1$ for $i\neq n_{\text{widget}}-1$.
@@
$\begin{split}t_{\text{hardware}}^{total}=t^{\text{total}}_{\text{intra}}+t^{\text{total}}_{\text{handover-inter}}+t_{\text{decode-delay}}^{\text{total}}.\end{split}$ (22)
In this equation, $t_{\text{decode-delay}}^{\text{total}}$ represents the overall delay involved in performing decoding, entirely using classical methods, for all QEC cycles. We propose a method to estimate $t^{\text{decode-delay}}_{tot}$ in section A.1, based on generic decoders.
@@
input : Look-up table: DistillationWidgets (Code, $C$, $Q$, $p_{\text{out}}$)
output : Total required physical qubits, $N_{\text{alloc-phys}}=4\tilde{n}_{\text{logical}}d^{2}+n^{\text{modules}}_{\text{per-leg}}(n_{\text{distilleries}}Q+2d^{2}l_{\text{tansfer-bus}})$
@@
4$\tilde{n}_{\text{logical}}$, $\tilde{S}_{\text{consump}}$, $\tilde{S}_{\text{prep}}$, $\tilde{S}_{\text{meas}}$} = Cabaliser-Tracker-Scheduler(QuantumCircuit);
@@
8            $n_{\text{distilleries}}\rightarrow 4n_{\text{distilleries}}$;
@@ 11             $p_{\text{log-cell}}=1-(1-\kappa(p/p_{\text{thresh}})^{(d+1)/2})^{d}$;
12             $M_{\varepsilon}=p_{\text{log-cell}}>\epsilon$;
@@ 14                  Decrease $\varepsilon$;
16      $M_{\text{log-cell}}=p_{\text{log-cell}}>p_{\text{out}}$;
@@ 18            Break loop;
Algorithm 1 RRE’s estimation methodology based on the middle-of-the-way architecture of fig. 2.
II Test cases
To demonstrate resource estimation using graph-state formalism, we generated circuits for quantum algorithms that target core computational challenges in real-world applications. Resource-efficient compilation necessarily means considering aspects of the hardware in the synthesis methodology and empirically evaluating alternatives when analytical results are not easily obtained. As such, these circuits are then compiled into fault-tolerant quantum computing patterns with specificity to superconducting architectures. This section summarizes the test cases chosen for circuit synthesis to demonstrate our resource estimation methodology.
II.1 Quantum Fourier Transform
The Quantum Fourier Transform (QFT) is a common component of quantum algorithms, including Quantum Phase Estimation. From the perspective of hardware-specific resource analysis, QFTs are likely to feature in application-relevant resource analysis, yet can be scaled down to a small handful of logical input gates, allowing us to trace and debug compilation workflows in our tool chain at the level of logical operations. This makes QFT circuits ideally suited as first test cases of our estimation framework. This work analyzes QFT circuits on $n$ qubits where $n=4,10,25,50,75,100,250,500$ or $1000$.
II.2 Hamiltonian simulation
Simulation of lattice-based Hamiltonians covers a family of critical applications investigated by researchers in quantum chemistry and condensed matter. Moreover, simulating dynamics may provide utility for smaller hardware system sizes than methods for evaluating static properties, such as the ground state.
Hamiltonian simulation consists of evolving an initial state |$\ket{\psi_{0}}$ to a state at time $t$, |$\ket{\psi}$ according to,
|$\ket{\psi}=e^{-iHt}\ket{\psi_{0}}$ (23)
where $H$ is the Hamiltonian of the system of interest.
We consider the simulation of two families of Hamiltonians: 2D Transverse-Ising models and 2D Fermi-Hubbard models. Each of these will be described in more detail below.
II.2.1 Transverse-Ising Hamiltonian Simulation
The Los Alamos National Laboratory (LANL) has created a public repository containing circuit generation and analysis of problems that may be amenable to quantum computing, and ostensibly are of interest to the scientific community, [72]. Drawing inspiration from their investigation of magnetic lattices, we perform resource estimation for the simulation of a time-invariant Transverse-Ising Hamiltonian,
$H=\sum_{\langle\textbf{jk}\rangle}\,Z_{\textbf{j}}\otimes Z_{\textbf{k}}+0.1\sum_{\textbf{j}}\,X_{\textbf{j}},$ (24)
on a triangular lattice, shown in Figure 8.
Figure 8: Triangular lattice of size $4\times 4$
@@ II.2.1 Transverse-Ising Hamiltonian Simulation
Based on the studies documented by LANL, we carried out resource estimates of simulation of the Hamiltonian in eq. 24 for a total time of $t=1$ (in the inverse Hamiltonian units) on a 2D triangular lattice of size $N\times N$ where $N$ is an integer between $2$ and $19$.
II.2.2 Fermi-Hubbard Hamiltonian simulations
We also investigated the Fermi-Hubbard simulation as a core computational capability for solving room-temperature superconductivity problems. These problems typically involve studying the ground state energy, state configurations, and other observable system properties, including time-dependent quantities.
@@ II.2.2 Fermi-Hubbard Hamiltonian simulations
where single and double angled brackets denote all nearest neighbor and next-nearest-neighbor lattice site pairs respectively. $c^{\dagger}_{\textbf{j},\sigma}(c_{\textbf{j},\sigma}$) creation (annihilation) operators of a spin-$\sigma$ Fermion on the jth lattice site where $\sigma\in\{\uparrow,\downarrow\}$. $J$ (primed or unprimed) and $U$ capture the magnitudes of the hopping and repulsion energies respectively. $n_{\textbf{k},\sigma}=c^{\dagger}_{\textbf{k},\sigma}c_{\textbf{k},\sigma}$ is the number operator, $h_{z}$ is an external applied magnetic field, and $\mu$ is the chemical potential.
@@
To motivate the specific Fermi-Hubbard simulations, we will obtain resource estimates. We will consult the literature to understand configurations that interest the community. Broadly speaking, methods for simulating a Fermi-Hubbard system can be broken into stochastic and deterministic families. As mentioned in [73], deterministic methods do not “suffer from numerical issues seen in stochastic simulations when dealing with systems with non-zero chemical potential”. Moreover, this instability due to the “sign problem” becomes prohibitive for moderate $\mu\sim 0.3$. This suggests separating instances into those with $\mu=0$, and those with $\mu>0.3$. Here, we reference a few exemplary studies in the literature on stochastic and deterministic methods.

### v3

@@ II.4.7 Space-time-optimal resource estimation
Here, we have defined the hybrid (intra- and inter-modular) time required to prepare the subgraph for preparation step $i$ as $t^{\text{prep}}_{i}=8d(N_{i}^{\text{intra-comms}}t+N_{i}^{\text{cross-module}}t_{\text{inter}})$. Where $N_{i}^{\text{intra-comms}}=\sum_{j\in L_{i,j}^{\text{prep}}}\left\lfloor\frac{\mathbb{L}_{i}^{\text{prep}}}{\sum_{k\in L_{i,j}^{\text{prep}}}\lfloor d^{\text{max}}_{i,j,k}/\tilde{n}_{\text{logical}}\rfloor}\right\rfloor~\text{for}~n^{\text{modules}}_{\text{per-leg}}>1$ is the count of intra-module operations, while $N_{i}^{\text{cross-module}}=\sum_{j\in L_{i}^{\text{prep}}}\left\lceil\frac{\sum_{k\in L_{i,j}^{\text{prep}}}\lfloor d^{\text{max}}_{i,j,k}/\tilde{n}_{\text{logical}}\rfloor}{n_{\text{inter-pipes}}}\right\rceil~\text{for}~n^{\text{modules}}_{\text{per-leg}}>1$ represents the number of cross-module operations in the preparation sub-schedule $i$ (only in the vertical direction along a ladder leg). We assume that $t_{\text{inter}}\gg t$ is the characteristic physical timescale for all inter-module communications. Moreover, we define $d^{\text{max}}_{i,j,k}$ as the maximum distance between any pairs of graph nodes in the tuple $k$ of sub-list $j$ of preparation sub-schedule $i$ (see [41] for details), and $n_{\text{inter-pipes}}$ is the number of inter-modular pipes connecting each pair of modules in the macro-architecture.
@@
$t^{\text{distill-delay}}_{\text{i}}=\begin{cases}8tC&\left\lceil\frac{L^{\text{max-seq}}_{\text{consump},i}-n^{T-\text{per-module}}_{i}}{n_{\text{T-factories}}}\right\rceil,\\ &\text{if }L^{\text{max-seq}}_{\text{consump},i}>n^{T-\text{per-module}}_{i}\\ 0,&\text{otherwise}\end{cases}$
@@
Similarly, for step $i$, an idle delay must be added if preparation for the next subgraph $i+1$ is not yet complete, and this delay runs in parallel across the interleaving modules in the other leg of the ladder. An upper bound on the individual time for intra-modular node consumption in step $i$ is $t^{\text{consump-intra}}_{i}=8td(\left\lceil\frac{n^{\text{max}-T}_{i}}{n_{\text{T-factories}}}\right\rceil+L_{\varepsilon}\left\lceil\frac{n^{\text{max}-Rz}_{i}}{n_{\text{T-factories}}}\right\rceil)$. Therefore, the preparation delay for step $i\in\{0,\dots,n_{\text{widgets}}-2\}$ becomes
@@
Lastly, we need to account for the total time required on the actual FTQC to stitch the output nodes of the graph at step $i$, located along one leg of the ladder, to the input nodes of the graph at step $i+1$, situated along the other leg of the ladder, for $i\neq n_{\text{widget}}-1$. We assume we can perform the hand-over process on-the-fly by teleporting logical Bell pairs across the modules in a horizontal direction, which involves adding logical $CNOT$ or $CZ$ gates (or, for smaller problems, some equivalent $SWAP$ operations). We must incur a cost if FTQC needs to teleport graph nodes vertically within the same leg (which is infeasible in $O(1)$ operations). Therefore, an upper bound for this quantity can be expressed as
$\begin{split}t^{\text{total}}_{\text{handover-inter}}=8t_{\text{inter}}d\sum_{i\in\{0,\dots,n_{\text{widget}}-2\}}\left\lceil\frac{n^{\text{cross-module-IO}}_{i}}{n_{\text{inter-pipes}}}\right\rceil,\end{split}$ (21)
where we define the total number of communications between cryostat modules needed to hand over output nodes of graph $i$ to $i+1$ as $n^{\text{cross-module-IO}}_{i}=\sum_{\tilde{o}\in S^{i}_{\text{output}},\tilde{i}\in S^{i+1}_{\text{input}}}n^{i}_{\tilde{o},\tilde{i}}$; here, $n^{i}_{\tilde{o},\tilde{i}}$ is the number of crossings necessary to hand over node $\tilde{o}$ of graph $i$ to node $\tilde{i}$ of graph $i+1$ for $i\neq n_{\text{widget}}-1$.
@@
$\begin{split}t_{\text{hardware}}^{\text{total}}=t^{\text{total}}_{\text{intra}}+t^{\text{total}}_{\text{handover-inter}}+t_{\text{decode-delay}}^{\text{total}}.\end{split}$ (22)
In this equation, $t_{\text{decode-delay}}^{\text{total}}$ represents the overall delay incurred by performing decoding, entirely using classical methods, across all QEC cycles. We propose a method to estimate $t^{\text{decode-delay}}_{tot}$ in appendix˜A, based on generic decoders.
@@
input : Look-up table: $T$-factories ($C$, $Q$, $p_{\text{out}}$)
output : Total allocated physical qubits, $n_{\text{alloc-phys}}=4\tilde{n}_{\text{logical}}d^{2}+n^{\text{modules}}_{\text{per-leg}}(n_{\text{T-factories}}Q+2d^{2}l_{\text{tansfer-bus}})$
@@
[…]
@@
8    $n^{\prime}\rightarrow 4n^{\prime}$;
@@ 11    $p_{\text{log-cell}}=1-(1-\kappa(p/p_{\text{thresh}})^{(d+1)/2})^{d}$;
12    $M_{\varepsilon}=p_{\text{logical}}>\epsilon$;
@@ 14        Decrease $\varepsilon$;
16 $M_{\text{logical}}=p_{\text{logical}}>p_{\text{out}}$;
@@ 18     Break loop;
Algorithm 1 RRE’s estimation methodology based on the fixed logical architectures of section˜II.
III Test algorithms
To demonstrate resource estimation using the graph-state formalism, we generated circuits for selected pedagogical quantum algorithms that still target core computational challenges in real-world applications. Resource-efficient compilation necessarily involves hardware considerations in the synthesis methodology and the empirical evaluation of alternatives when analytical results are not readily available. As such, these circuits are then compiled into FT quantum computing patterns tailored to superconducting architectures presented in section˜II. This section summarizes the test cases selected for circuit synthesis to demonstrate our resource-estimation methodology.
III.1 Quantum Fourier Transform
The Quantum Fourier Transform (QFT) is a common component of quantum algorithms, including Quantum Phase Estimation. From the perspective of hardware-specific resource analysis, QFTs are likely to feature in application-relevant resource analysis, yet can be scaled down to a small handful of logical input gates, allowing us to trace and debug compilation workflows in our tool chain at the level of logical operations. This makes QFT circuits ideally suited as the first test cases of our estimation framework. This work analyzes QFT circuits on $n_{\text{input}}$ logical qubits, where $n_{\text{input}}=4,10,25,50,75,100,250,500$ or $1000$.
III.2 Hamiltonian simulations
Simulation of lattice-based Hamiltonians encompasses a family of critical applications investigated by researchers in quantum chemistry and condensed matter physics. Moreover, simulating dynamics may be more useful for smaller hardware systems than methods that evaluate static properties, such as the ground state.
[…]
where $H$ is the Hamiltonian of the system of interest. We consider the simulation of two families of Hamiltonians: 2D Transverse-Ising models and 2D Fermi-Hubbard models. Each of these will be described in more detail below.
III.2.1 Transverse-Ising Hamiltonian Simulations
The Los Alamos National Laboratory (LANL) has created a public repository containing circuit generation and problem analysis for problems that may be amenable to quantum computing and are of interest to the scientific community [54]. Drawing inspiration from their investigation of magnetic lattices, we perform resource estimation for the simulation of a time-invariant Transverse-Ising Hamiltonian,
$H=\sum_{\langle\textbf{jk}\rangle}\,Z_{\textbf{j}}\otimes Z_{\textbf{k}}+0.1\sum_{\textbf{j}}\,X_{\textbf{j}}.$ (24)
@@ III.2.1 Transverse-Ising Hamiltonian Simulations
Based on the studies documented by LANL, we carried out resource estimates of simulation of the Hamiltonian in eq.˜24 for a total time of $t=1$ (in the inverse Hamiltonian units) on a 2D triangular lattice (sometimes called a 2D hexagonal Bravais lattice) of size $N\times N$ where $N$ is an integer between $2$ and $19$.
III.2.2 Fermi-Hubbard Hamiltonian Simulations
We also investigated the Fermi-Hubbard simulation as a core computational capability for solving problems in room-temperature superconductivity [3]. These problems typically involve studying the ground state energy, state configurations, and other observable system properties, including time-dependent quantities.
@@ III.2.2 Fermi-Hubbard Hamiltonian Simulations
where single and double angled brackets denote all nearest-neighbor and next-nearest-neighbor lattice site pairs, respectively. $c^{\dagger}_{\textbf{j},\sigma}(c_{\textbf{j},\sigma}$) creation (annihilation) operators of a spin-$\sigma$ fermion on the jth lattice site where $\sigma\in\{\uparrow,\downarrow\}$. $J$ (primed or unprimed) and $U$ capture the magnitudes of the hopping and repulsion energies, respectively. $n_{\textbf{k},\sigma}=c^{\dagger}_{\textbf{k},\sigma}c_{\textbf{k},\sigma}$ is the number operator, $h_{z}$ is an external applied magnetic field, and $\mu$ is the chemical potential.
@@
To motivate the specific Fermi-Hubbard simulations for which we will obtain resource estimates, we reviewed the literature to identify configurations of interest to the community. Broadly speaking, methods for simulating a Fermi-Hubbard system can be classified as stochastic or deterministic. As mentioned in [79], deterministic methods do not “suffer from numerical issues seen in stochastic simulations when dealing with systems with non-zero chemical potential”. Moreover, this instability due to the “sign problem” becomes prohibitive for moderate $\mu\sim 0.3$. This suggests separating instances into those with $\mu=0$ and those with $\mu>0.3$. Here, we reference a few exemplary studies in the literature on stochastic and deterministic methods.
Some examples of stochastic studies include a study using Auxiliary-field quantum Monte Carlo (AFQMC) to estimate ground states on lattice sizes of $4\times 4$ to $16\times 16$, $U/J=4,8$, and $J^{\prime}=0$. [66]. In addition, [31] used Determinant Quantum Monte Carlo (DQMC) to study doped ground state estimation using $U/J=6$, $J^{\prime}/J=-.25$, and lattice sizes of $8\times 8$, $\times 8, 12×12$$\times 12$, $16\times 8$. In contrast, examples of deterministic studies include work on numerical linked-cluster expansions [65]. In this study, researchers selected random repulsion strengths from an uncertainty “box” about $U_{0}$. Simulations were for a periodic $10\times 10$ lattice with $U_{0}/J=8$, $J^{\prime}=0$. Other researchers used Projected Density Matrix Embedding (PDME) to perform a 2D Fermi-Hubbard study with $U/J=2,4,6,8$, $J^{\prime}=0$ on a $40\times 40$ lattice [89]. Finally, in [79], the authors use Projected Entangled Pair States (PEPS) with a bond dimension of $D=10$, and claim that accuracy increases with $D$. Most of their results are from a $3\times 4$ honeycomb lattice with $U/J\in[0,6]$ and $J^{\prime}=0$. Still, the authors were able to calculate the ground state via imaginary time evolution on a $15\times 15$ bipartite honeycomb lattice.
Of note is that in [79] it was determined that the time required to run these studies on a CPU varies like, $t_{\text{C}PU}\propto L_{x}L_{y}(2\chi^{3}D^{4}d^{2}+\chi^{2}D^{2}d^{2}+\chi^{2}D^{4}d^{4}),$ where $\chi$ is the truncated bond dimension and $d$ the problem dimension. Memory scaled as, $2\chi^{2}D^{4}d^{2}M_{\text{n}um}+2N\chi^{2}D^{2}dM_{\text{n}um}$. The authors empirically found that $\chi=3D$ was sufficiently accurate for their studies, with an error of about $1\%$. These estimates provide a useful guidepost for the tipping point at which FTQCs may outperform classical computers on these types of problems.
III.2.3 Quantum algorithm and approaches for Hamiltonian simulations
Owing to its optimal scaling characteristics for queries to the block encoding of the Hamiltonian [43], we have chosen Quantum Signal Processing (QSP) as the Hamiltonian simulation algorithm for which we obtain a resource estimate. QSP is briefly described below, but more details can be found in [43], [44], [51], and [12].
Given a Hamiltonian $H$ and simulation time $t$, QSP performs a sequence of single-qubit rotations, parameterized by a sequence of phase angles, interlaced with a quantum walk-like operator. This allows one to apply a polynomial approximation of $e^{-iHt}$ to the input target state. The sequence of phase angles determines the polynomial approximation. It is computed offline, using a classical algorithm, with the desired error tolerance specified by the user-defined parameter $\epsilon_{\text{qsp}}$. $\epsilon_{\text{qsp}}$ is also inversely related to the probability of success of the QSP algorithm. Thus, for an accurate simulation, with a high probability of success, $\epsilon_{\text{qsp}}$ may have to be set relatively small. For implementation details, we follow the examples accompanying the pyLIQTR [59] package v1.3.3, as well as the work done in [50, 93].
For demonstration purposes, we set the desired probability of success of our algorithm, $p_{\rm algo}$, to be equal to 0.95. In real-world applications, this is typically fixed to meet outcome requirements (e.g., accuracy and wall-time). Denoting the probability of success of a QSP circuit, $p_{\rm qsp}$, we therefore require,
$p_{\rm algo-fail}\leq p_{\rm qsp}.$ (26)
Since circuit size is inversely related to $\epsilon_{\text{qsp}}$, it is prudent to choose $\epsilon_{\text{qsp}}$ to be as large as possible while still guaranteeing that $p_{\rm qsp}$ exceeds the algorithm target of 0.95. According to [43] $(1-2\epsilon_{\text{qsp}})\leq p_{\rm qsp}$. Thus, to ensure eq.˜26 is satisfied we set,
$1-2\epsilon_{\text{qsp}}=p_{\rm algo-fail},$ (27)
and so in this work we choose $\epsilon_{\text{qsp}}$ to be equal to $0.025$.
III.3 Simulation circuit synthesis
For our Hamiltonian simulation circuits, we used pyLIQTR v1.3.3 to construct the Hamiltonians and synthesize the QSP circuits. The pyLIQTR package provides tools that construct the Hamiltonians, calculate the necessary phase angles for QSP simulation, and generate QSP circuits. In addition, pyLIQTR allows the user to request random phase angles rather than solving for the precise values required for an accurate simulation. In our work, we used random phase angles. This is appropriate because we only perform resource estimation, not numerically accurate simulations. The correct number of phase angles is used in the circuits, but their numerical values are random. This provides a conservative overestimate of the resource estimations without explicitly incurring the time-consuming step of solving for the phase angles. For Fermi-Hubbard models, we utilized pyLIQTR’s FermiHubbard model class, and employed the FermiHubbardSquare block encoding, which is based on [6].
IV Resource estimation results
Table 2: Rigetti’s resource estimation results on a superconducting fault-tolerant quantum computing architecture, section˜II, with space-optimized graph-state compilation. Here, “# Logical bus patches” includes patches for the memory and auxiliary buses that mediate logical measurements in the bilinear pattern shown in fig.˜1. Likewise, “# $T$ patches” includes logical patches for all $T$-factories and $T$-transfer Buses.
Input circuit Logical resources Superconducting FTQC resources
Instance $\tilde{n}_{\text{logical}}$ # Logical bus patches # T patches Total allocated phys. qubits, $n_{\text{alloc-phys}}$ Total modules, $2n^{\text{modules}}_{\text{per-leg}}$ $t_{\text{tot}}^{\text{hardware}}$ $P_{\text{Diss},\text{4K}}$ $P_{\text{Diss},\text{20mK}}$ $E_{\text{tot}}$
Fermi Hubbard
2x2 36 280 4.05k 1.32M 2 4.09s 840W 168nW 668Wh
4x4 74 332 3.32k 1.21M 2 35.1s 840W 168nW 5.73kWh
8x8 182 728 2.07k 1.38M 2 9.49m 840W 168nW 93kWh
16x16 578 2.31k 1.25k 3.30M 4 6.9h 1.68kW 336nW 8.11MWh
20x20 878 3.52k 1.4k 4.96M 6 37.8h 2.52kW 504nW 66.7MWh